# Exploring How Audio Effects Alter Emotion with Foundation Models

### v1

@@ 1 Introduction
At the same time, advances in foundation models—large-scale neural architectures pre-trained on multimodal datasets, have demonstrated an ability to capture high-level semantic and perceptual relationships across domains, including with audio FX [10, 11]. In the music domain, foundation models such as MusicLM [12], MERT [13], Qwen [14], and CLAP [15] have shown promising abilities in generating emotionally expressive audio, learning disentangled latent representations, and enabling cross-modal retrieval tasks. These models implicitly encode rich associations between musical structure, timbre, and affective meaning, making them ideal tools to probe novel hypotheses about music-emotion relationships.
@@ 3.4 Real World Scenario
As shown in Fig. 3, moving from isolated effects to complete FX chains highlights how cumulative production choices shift embeddings in latent space. MERT and Qwen show similar, structured trajectories reflecting these cumulative effects. The distortion-heavy chain of Rage Against the Machine produces nearly unidirectional shifts, while the reverb- and delay-heavy chains of U2 and Pink Floyd show more complex, diffuse paths. CLAP exhibits shorter, scattered movements, indicating reduced sensitivity to combined effects. Overall, real-world FX chains induce larger, more coherent shifts, suggesting that trajectory patterns—unidirectional versus multidirectional—can serve as proxies for how effect categories shape emotional perception.
Fig. 3: UMAP visualization of foundation model embeddings for the witheFlow dataset, showing trajectories generated for each input after applying audio FX with the intensity ranging from 1 to 10.

### v2

@@ 1 Introduction
At the same time, advances in foundation models—large-scale neural architectures pre-trained on multimodal datasets, have demonstrated an ability to capture high-level semantic and perceptual relationships across domains, including with audio FX [10, 11]. In the music domain, foundation models such as MusicLM [12], MERT [13], Qwen [14], and CLAP [15] have shown promising abilities in generating emotionally expressive audio, learning disentangled latent representations, and enabling cross-modal retrieval tasks [16]. These models implicitly encode rich associations between musical structure, timbre, and affective meaning, making them ideal tools to probe novel hypotheses about music-emotion relationships.
@@ 3.4 Real World Scenario
As shown in Fig. 3, moving from isolated effects to complete FX chains reveals how cumulative production choices shift embeddings in latent space. MERT and Qwen produce similar trajectories, both showing structured, directional shifts that reflect the cumulative impact of effects. The distortion-heavy chain of Rage Against the Machine yields almost unidirectional trajectories, consistent with strong spectral shaping imposing a uniform transformation. Reverb- and delay-heavy chains of U2 and Pink Floyd also follow a generally consistent pattern, though some divergent paths reflect the more complex ways spatial and temporal effects influence latent representations. CLAP, in contrast, shows shorter, scattered movements, suggesting that its sensitivity is dampened when effects are combined. Overall, real-world effect chains generate larger, more coherent shifts, highlighting that artists naturally select combinations to evoke stronger emotional responses. The nature of these trajectories—unidirectional versus multidirectional—thus serves as a proxy for how different categories of audio effects shape emotional perception in foundation models.
Fig. 3: UMAP visualization of foundation model embeddings for the witheFlow dataset, showing trajectories generated for each input after applying real-world scenario audio FX.